// Elo ratings 547,244 matches rated · 14,934 teams
Strength, and what it says about draws

How strong is a team, and how often do equals draw?

Every match in the database is replayed in date order. Both sides start at 1500; the winner takes points from the loser, more when the result was a surprise and more again when the margin was wide. A team's rating is therefore what its results have earned, and the gap between two ratings before kick-off is a measure of how uneven the match should be. The tables below show how often each size of gap ended level.

Matches rated547,244
Teams rated14,934
Evenly matched draw 27.37% gap under 25
Mismatch draw 13.95% gap 300+
Descriptive statistics, not betting advice → read this
Clubs: how often each rating gap drew 497,563 matches
Clubs: how often each rating gap drew 0% 15% 30% drew 0–24 — bookmaker's figure 28.79% 0–24 — drew 27.37% 27.4% 0–24 25–49 — bookmaker's figure 28.66% 25–49 — drew 26.30% 26.3% 25–49 50–74 — bookmaker's figure 28.34% 50–74 — drew 25.02% 25.0% 50–74 75–99 — bookmaker's figure 27.91% 75–99 — drew 25.20% 25.2% 75–99 100–149 — bookmaker's figure 26.93% 100–149 — drew 25.26% 25.3% 100–149 150–199 — bookmaker's figure 25.26% 150–199 — drew 23.71% 23.7% 150–199 200–299 — bookmaker's figure 22.31% 200–299 — drew 20.56% 20.6% 200–299 300+ — bookmaker's figure 16.00% 300+ — drew 13.95% 13.9% 300+ rating gap before kick-off
drew bookmaker's figure How often each rating gap actually drew, beside the draw chance the closing odds implied.
Rating gap Matches Drew Bookmaker's figure Return if you bet every draw
0–24 67,464 27.37% 28.79% -1.08%
25–49 73,116 26.30% 28.66% -3.14%
50–74 79,313 25.02% 28.34% -2.77%
75–99 64,945 25.20% 27.91% -2.60%
100–149 87,110 25.26% 26.93% -3.11%
150–199 53,387 23.71% 25.26% -2.52%
200–299 49,646 20.56% 22.31% -5.79%
300+ 22,582 13.95% 16.00% -12.49%

Evenly matched sides draw far more often than mismatched ones — the first row against the last is the whole story. The last three columns compare that with the betting market on the matches that carry a price: the bookmaker's figure tracks reality closely, and betting every draw loses money in every row. "Drew" and "Drew (those)" differ because priced matches are top divisions, which draw slightly more often.

National teams: how often each rating gap drew 49,681 matches
National teams: how often each rating gap drew 0% 15% 30% drew 0–24 — drew 27.50% 27.5% 0–24 25–49 — drew 25.37% 25.4% 25–49 50–74 — drew 26.90% 26.9% 50–74 75–99 — drew 25.86% 25.9% 75–99 100–149 — drew 24.58% 24.6% 100–149 150–199 — drew 23.11% 23.1% 150–199 200–299 — drew 19.92% 19.9% 200–299 300+ — drew 12.37% 12.4% 300+ rating gap before kick-off
drew How often each rating gap actually drew.
Rating gap Matches Drew Bookmaker's figure Return if you bet every draw
0–24 5,179 27.50% — —
25–49 5,077 25.37% — —
50–74 5,115 26.90% — —
75–99 4,618 25.86% — —
100–149 8,299 24.58% — —
150–199 6,447 23.11% — —
200–299 8,423 19.92% — —
300+ 6,523 12.37% — —
Top clubs by current rating
# Team Rating Peak
1 FC Bayern München 2,301 2,328
2 Arsenal 2,255 2,293
3 Manchester City 2,213 2,331
4 Paris Saint-Germain 2,212 2,231
5 Barcelona 2,203 2,243
6 Real Madrid 2,178 2,237
7 Liverpool 2,121 2,285
8 Brighton & Hove Albion 2,120 2,123
9 Bournemouth 2,114 2,114
10 FK Bodø/Glimt 2,112 2,143
11 Inter Milan 2,108 2,202
12 Aston Villa 2,102 2,162
13 Borussia Dortmund 2,085 2,116
14 Atlético Madrid 2,079 2,134
15 Viking FK 2,076 2,097
16 SL Benfica 2,051 2,071
17 Brentford 2,051 2,070
18 Chelsea 2,049 2,185
19 Bayer Leverkusen 2,046 2,190
20 Sporting CP 2,044 2,145
21 AS Roma 2,033 2,070
22 VfB Stuttgart 2,027 2,050
23 Como 1907 2,024 2,024
24 Club Brugge KV 2,016 2,023
25 Leeds United 2,015 2,033
26 Zenit Saint Petersburg 2,010 2,011
27 Royale Union Saint-Gilloise 2,009 2,037
28 Real Betis 1,996 1,999
29 Everton 1,995 2,009
30 Porto 1,995 2,086
Top national teams by current rating
# Team Rating Peak
1 Spain 2,312 2,312
2 England 2,195 2,195
3 Argentina 2,190 2,200
4 France 2,157 2,210
5 Portugal 2,087 2,088
6 Morocco 2,075 2,106
7 Brazil 2,039 2,212
8 Colombia 2,033 2,115
9 Switzerland 2,033 2,033
10 Japan 2,026 2,064
11 Netherlands 2,023 2,177
12 Belgium 2,018 2,175
13 Norway 1,973 2,053
14 Germany 1,965 2,254
15 Mexico 1,965 2,054
16 Italy 1,957 2,095
17 Nigeria 1,949 1,962
18 Australia 1,936 2,002
19 Senegal 1,924 1,991
20 Croatia 1,923 2,036
21 United States 1,923 2,011
22 Ecuador 1,919 1,980
23 Austria 1,917 1,957
24 Algeria 1,911 1,961
25 Uruguay 1,909 2,056
26 Denmark 1,909 2,018
27 Paraguay 1,905 1,905
28 South Korea 1,901 1,986
29 Iran 1,900 2,017
30 Canada 1,878 2,010
All rated teams 14,934 teams
Clear
▸ Reset
# Team Rating Peak
1 Spain 2,312 2,312
2 FC Bayern München 2,301 2,328
3 Arsenal 2,255 2,293
4 Manchester City 2,213 2,331
5 Paris Saint-Germain 2,212 2,231
6 Barcelona 2,203 2,243
7 England 2,195 2,195
8 Argentina 2,190 2,200
9 Real Madrid 2,178 2,237
10 France 2,157 2,210
11 Liverpool 2,121 2,285
12 Brighton & Hove Albion 2,120 2,123
13 Bournemouth 2,114 2,114
14 FK Bodø/Glimt 2,112 2,143
15 Inter Milan 2,108 2,202
16 Aston Villa 2,102 2,162
17 Portugal 2,087 2,088
18 Borussia Dortmund 2,085 2,116
19 Atlético Madrid 2,079 2,134
20 Viking FK 2,076 2,097
21 Morocco 2,075 2,106
22 SL Benfica 2,051 2,071
23 Brentford 2,051 2,070
24 Chelsea 2,049 2,185
25 Bayer Leverkusen 2,046 2,190
26 Sporting CP 2,044 2,145
27 Brazil 2,039 2,212
28 Colombia 2,033 2,115
29 Switzerland 2,033 2,033
30 AS Roma 2,033 2,070
31 VfB Stuttgart 2,027 2,050
32 Japan 2,026 2,064
33 Como 1907 2,024 2,024
34 Netherlands 2,023 2,177
35 Belgium 2,018 2,175
36 Club Brugge KV 2,016 2,023
37 Leeds United 2,015 2,033
38 Zenit Saint Petersburg 2,010 2,011
39 Royale Union Saint-Gilloise 2,009 2,037
40 Real Betis 1,996 1,999
41 Everton 1,995 2,009
42 Porto 1,995 2,086
43 Manchester United 1,995 2,147
44 Rangers 1,990 2,087
45 Tottenham Hotspur 1,989 2,136
46 RB Leipzig 1,989 2,145
47 Newcastle United 1,987 2,141
48 PSV Eindhoven 1,980 2,089
49 AC Milan 1,976 2,071
50 Norway 1,973 2,053
page 1 / 299 ‹ prev next ›
Reading betting odds why the three prices add up to more than 100%

The prices on the fixtures page are decimal odds: what one unit returns, your stake included. At 3.50, a 20 Kč bet on the draw returns 70 Kč — 50 profit plus the 20 back.

Turn a price into a probability by dividing it into 1. 1 / 3.50 = 28.6%: that is how likely the bookmaker's price says a draw is. Do it for all three results of a match and they should add up to 100%. They never do.

The excess is the bookmaker's margin — its built-in profit, also called the overround. Spread the excess back out and you get the "fair" figures used in the tables above.

This is why betting every draw in any row above loses money: the prices are already close to what really happens, minus the margin. To profit you need odds longer than 1 / your own probability — on a team that draws a quarter of the time, better than 4.00.

Result Odds 1 / odds Margin removed
Home win 2.00 50.0% 48.3%
Draw 3.50 28.6% 27.6%
Away win 4.00 25.0% 24.1%
Total — 103.6% 100.0%

The three prices imply 103.6% of probability between them, so about 3.6% belongs to the bookmaker whatever happens. That cut is taken on every bet, win or lose, which is why the "betting them all" column above is negative even where the odds look generous.

FDB4 rates every team with Elo, replayed in date order over more than half a million matches: both sides start at 1500, the winner takes points from the loser, and a surprise or a wide margin moves the rating further. The page ranks the strongest clubs and national teams by current rating and peak, then shows how often each size of rating gap actually ended level — evenly matched sides draw roughly twice as often as mismatched ones. Alongside it sits the betting market's own figure for the same matches and a plain explanation of how decimal odds, implied probability and the bookmaker's margin fit together.